OpenStudy (anonymous):

which of the following are solutions to the equation x^2+5x+3=9 a-5 b3 c1 d-3 e-6 f-1

OpenStudy (akashdeepdeb):

Here goes x^2+5x+3=9 x^2+5x-6=0 x^2+(6-1)x-6=0 x^2 + 6x - x -6 =0 x(x+6) - 1(x+6) = 0 (x-1)(x+6) = 0 Thus x can have values 1,-6 c AND e are the solutions :) I hope this helped, Kudos Plz? :D

OpenStudy (anonymous):

lol